OpenStudy (anonymous):

What is the equation of the line that has a slope of –2 and a y-intercept of 0? A. y = 2 B. y = –2 C. y = 0 D. y = –2x

OpenStudy (rsadhvika):

y = mx + c

OpenStudy (rsadhvika):

m = -2 c = 0 plugin

OpenStudy (rsadhvika):

wat do you get ? :)

OpenStudy (anonymous):

i got same thing just wanted to be sure

OpenStudy (rsadhvika):

which one ?

OpenStudy (anonymous):

C

OpenStudy (rsadhvika):

omg ! no

OpenStudy (anonymous):

????

OpenStudy (rsadhvika):

y = mx + c put m = -2 c = 0 y = -2x + 0 y = -2x

OpenStudy (anonymous):

oh okay im not good at math thank you :)

OpenStudy (rsadhvika):

np :) it just takes practice

OpenStudy (anonymous):

yeah i can see that lol

OpenStudy (rsadhvika):

:D so the answer is D okay

OpenStudy (anonymous):

okay :) thanks again

OpenStudy (rsadhvika):

yw !